(x)=x^(2)-18 and G(x)=2-x
find (f.G)(n)

Answer by greenestamps(13200)



G(x) = 2-x --> G(n) = 2-n

(f.G)(n) = f(2-n) = (2-n)^2-18

Expand and simplify